OpenStudy (anonymous):

In Belinda's hometown, the sales tax rate is 7.5%. Belinda is shopping for a new dresser and finds one at a local furniture store for $672.00 plus sales tax. She finds the same dresser on the Internet for $550 from a furniture store in Maine. If she orders the dresser from Maine, there is an additional $95 shipping charge for 3-day delivery, and she will still have to pay the 7.5% sales tax on the dresser, but not on the delivery charge. How much money would Belinda save by ordering the dresser over the Internet instead of buying it locally?

OpenStudy (anonymous):

Locally it would cost her : 672 + 7.5% sales tax (672)(0.075) = 50.40 672 + 50.40 = 722.40 (local cost) Internet cost is : (550 + 7.5%) + 95 for shipping (550)(0.075) = 41.25 550 + 41.25 = 591.25 591.25 + 95 = 686.25 (internet cost) 722.40 - 686.25 = 36.15 Belinda would save $ 36.15 when ordering it through the internet. Do you understand what I have just done ?